if sin x + cos y=2, then sin[(2x+y)/3] = ?

Grade:12th pass

1 Answers

Vijay Mukati
8 years ago
Sinx + cosy=2, possible only if sinx =1 and cosy=1. That means, x = pi/2 and y=0. Therefore the value of sin[(2x+y)/3] = sin(60) = sqrt(3)/2.